OpenStudy (fibonaccichick666):

proof question, I need to prove that \[f'(c)=\lim_{h\rightarrow 0} \frac{f(c+h)-f(c-h)}{2h}\]

OpenStudy (fibonaccichick666):

So my issue is getting 2h on the bottom. Through a substitution from the definition of derivative of h=x-c, I can get the top and bottom minus the 2. Any ideas?

OpenStudy (fibonaccichick666):

@satellite73 any thoughts?

OpenStudy (freckles):

\[\lim_{h\rightarrow 0} \frac{f(c+h)-f(c-h)}{2h}\] \[\lim_{h \rightarrow 0}\frac{f(c+h)-f(c)}{2h}-\lim_{h \rightarrow 0}\frac{f(c-h)-f(c)}{2h}\]

OpenStudy (freckles):

then I think it should be good from here right?

OpenStudy (freckles):

pull out the constant multiple and you know f'(c) equals you know what

OpenStudy (fibonaccichick666):

ahh add c subtract c. Thanks

OpenStudy (fibonaccichick666):

oh wait, no that doesn't get me the 1/2

OpenStudy (freckles):

add f(c) and subtract f(c) and if you aren't convinced that one thing is f'(c) you could substitute the -h for u

OpenStudy (fibonaccichick666):

we cannot assume the conclusion

OpenStudy (freckles):

\[\frac{1}{2}f'(c)-\frac{1}{2} f'(c)\]

OpenStudy (freckles):

oh your right it gives you 0

OpenStudy (freckles):

Let f(x)=x^2 f(c)=c^2 f(c+h)=(c+h)^2=c^2+2ch+h^2 f(c-h)=(c-h)^2=c^2-2ch+h^2\[\lim_{h \rightarrow 0}\frac{c^2+2ch+h^2-c^2+2ch-h^2}{2h}=\lim_{h \rightarrow 0}\frac{2ch+2ch}{2h}=\lim_{h \rightarrow 0}(c+c)=2c\] lol just had to make sure

OpenStudy (fibonaccichick666):

I just cannot figure out the two... but if I compare it to the definition I know that involves h \[f'(x)=\lim_{h\rightarrow 0}\frac{f(x+h)-f(x)}{h}\]We haven't proved that, but it is the def I know from calc... so somehow, f(x)=1/2f(x-h)

OpenStudy (fibonaccichick666):

I get that lol

OpenStudy (fibonaccichick666):

I didn't buy it at first either

OpenStudy (fibonaccichick666):

I am trying to turn \[\lim_{x\rightarrow c}\frac{f(x)-f(c)}{x-c}\] into that

OpenStudy (freckles):

yeah that is what i'm trying to do now

OpenStudy (fibonaccichick666):

So my first step was a substitution of h=x-c which turns it into \[\lim_{h\rightarrow 0}\frac{f(c+h)-f(x-h)}{h}\] Ahhh! we have to change the x-h into a c-h right!

OpenStudy (freckles):

\[\lim_{u \rightarrow c}\frac{f(u+2h)-f(u)}{2(c-u)}\] hmmm...

OpenStudy (fibonaccichick666):

so we input so we put in the c+h for the x... no that doesn't work. We just get c then

OpenStudy (fibonaccichick666):

wouldn't yours have to be a double limit?

OpenStudy (freckles):

I think I made a typeo let's see u=c-h u+h=c h=c-u so c+h=c+(c-u)=2c-u

OpenStudy (freckles):

\[\lim_{u \rightarrow c}\frac{f(2c-u)-f(u)}{2(c-u)}\]\[\lim_{u \rightarrow c}\frac{f(2c-u)-f(u)}{2c-2u}\]

OpenStudy (freckles):

Like the bottom would have been cooler if we had (2c-u)-u but we don't

OpenStudy (fibonaccichick666):

yea... hmm

OpenStudy (fibonaccichick666):

why do you have u=c-h?

OpenStudy (freckles):

was making a substitution to see if i could get it in the form\[\lim_{u \rightarrow c}\frac{f(u)-f(c)}{u-c}\]

OpenStudy (freckles):

or the other way around whatever \[\lim_{u \rightarrow c}\frac{f(c)-f(u)}{c-u}\]

OpenStudy (freckles):

\[f'(c)=\lim_{h \rightarrow 0}\frac{f(c+h)-f(c)}{h} \\ f'(c)=\lim_{-h \rightarrow 0}\frac{f(c-h)-f(c)}{-h}=\lim_{h \rightarrow 0}\frac{f(c)-f(c-h)}{h} \\ f'(c)+f'(c)=\lim_{h \rightarrow 0}\frac{f(c+h)-f(c)}{h}+\lim_{h \rightarrow 0}\frac{f(c)-f(c-h)}{h}\]

OpenStudy (fibonaccichick666):

we can't go from that def because we haven't proved it :/

OpenStudy (freckles):

but the other one the one mentioned just a second ago?

OpenStudy (fibonaccichick666):

I did, we used it in calc, but since we haven't proved it in this class, we are not allowed to use it without deriving it

OpenStudy (freckles):

\[\lim_{u \rightarrow c}\frac{f(u)-f(c)}{u-c}=f'(c)?\]

OpenStudy (freckles):

can we use that one?

OpenStudy (fibonaccichick666):

what was your u again? or is it the same as x?

OpenStudy (fibonaccichick666):

if it is the same as x,, yes we can use it

OpenStudy (freckles):

yeah you can think of u as x it is just a variable

OpenStudy (fibonaccichick666):

you had a substitution prior, so I wasn't sure if you were still using it

OpenStudy (freckles):

I will use x if you need me to \[\lim_{x \rightarrow c}\frac{f(x)-f(c)}{x-c}=f'(c)\] let x-c=h so we have as x->c then h->0 \[\lim_{h \rightarrow 0}\frac{f(c+h)-f(c)}{h}=f'(c)\] you just proved it now you can use it

OpenStudy (fibonaccichick666):

meh, ok, that's what I showed earlier, only issue was the secondary x substitution on f(x-h) I was thinking it may not be alright

OpenStudy (freckles):

What? I don't understand?

OpenStudy (freckles):

We just showed that following are equivalent: \[\lim_{x \rightarrow c}\frac{f(x)-f(c)}{x-c}=\lim_{h \rightarrow 0}\frac{f(c+h)-f(c)}{h}=f'(c)\] \[f'(c)=\lim_{h \rightarrow 0}\frac{f(c+h)-f(c)}{h} \\ f'(c)=\lim_{-h \rightarrow 0}\frac{f(c-h)-f(c)}{-h}=\lim_{h \rightarrow 0}\frac{f(c)-f(c-h)}{h} \\ f'(c)+f'(c)=\lim_{h \rightarrow 0}\frac{f(c+h)-f(c)}{h}+\lim_{h \rightarrow 0}\frac{f(c)-f(c-h)}{h} \]

OpenStudy (fibonaccichick666):

well in order to get the f(c) you have to substitute in x=h+c

OpenStudy (fibonaccichick666):

that was where I was having an algebra moment of forgetfulness, but anyways, I still don't see how we get 2h alone on the bottom

OpenStudy (freckles):

what is f'(c)+f'(c)

OpenStudy (fibonaccichick666):

...ah ok I missed that equality

OpenStudy (fibonaccichick666):

ok, I got it from here, I just have to add justification to the left limit and hope they accept it. I still do not know how to properly write up a left/right limit, since they technically don't exist in this class

OpenStudy (fibonaccichick666):

thanks again

OpenStudy (freckles):

what do you mean write up a left/right limit?

OpenStudy (freckles):

is that a different question?

OpenStudy (fibonaccichick666):

no, but I have to justify the left handed limit concept used to switch the top around.

OpenStudy (freckles):

you need to use left and right limits to prove this?

OpenStudy (freckles):

I'm sorry I'm having trouble understanding.

OpenStudy (fibonaccichick666):

You used the fact that the left hand limit is equal to the limit in your step 3

OpenStudy (freckles):

oh i didn't do anything with left and right limits

OpenStudy (fibonaccichick666):

you used -h, and h

OpenStudy (fibonaccichick666):

orrrrrrr those are two separate sequences. I can spin that

OpenStudy (freckles):

\[f'(c)=\lim_{h \rightarrow 0}\frac{f(c+h)-f(c)}{h} \\ h=-a \\ h->0 \text{ then } -a->0 \\ f'(c)=\lim_{-a \rightarrow 0}\frac{f(c-a)-f(c)}{-a} \ \text{ replace }a \text{ with } h \] but if -a->0 then a->0 \[f'(c)=\lim_{a \rightarrow 0}\frac{f(c-a)-f(c)}{-a}\]

OpenStudy (freckles):

but you can replace the a back with h

OpenStudy (freckles):

also people i think usually symbolize left limit like this h->0^-

OpenStudy (fibonaccichick666):

that is using the concept of left/right limits, but it's cool. Conceptually that is what is occurring. We can't say that \(lim{~a\rightarrow 0}=lim {-a \rightarrow 0}\)

OpenStudy (fibonaccichick666):

but anyways thanks, I got this question from here

OpenStudy (freckles):

\[\lim_{x \rightarrow 5}(x+6)=5+6=11 \\ \lim_{-x \rightarrow 5}(-x+6)=-(-5)+6=5+6=11\]

OpenStudy (freckles):

if a->0 then so does -a

OpenStudy (freckles):

it is like saying if a=0 then -a=0

OpenStudy (fibonaccichick666):

that's not the same question

OpenStudy (freckles):

i didn't use left and right derivatives or limits

OpenStudy (fibonaccichick666):

for your example(the second is \(-x\rightarrow -5\), and second. You have to rigorously prove it, we can't just use it

OpenStudy (freckles):

no i have -x->5

OpenStudy (fibonaccichick666):

but anyways, it doesn't matter we are arguing over nothing at the moment

OpenStudy (freckles):

you have to use substitutions to prove this

OpenStudy (fibonaccichick666):

then your limt would =1

OpenStudy (fibonaccichick666):

sorry, just your work was weird

OpenStudy (fibonaccichick666):

you shouldn't have -(-5) it should just be 5 based off of your limit

OpenStudy (freckles):

\[\lim_{x \rightarrow 5}(x+6)=5+6=11 \\ \lim_{-x \rightarrow 5}(-x+6)=-(-5)+6=5+6=11 \] Ok we have \[\lim_{x \rightarrow 5}(x+6) \text{ \let } x=-u \\ \text{ so we have } \lim_{-u \rightarrow 5}(-u+6)\]

OpenStudy (fibonaccichick666):

though thy are logically equivalent

OpenStudy (freckles):

so like I said those are both equal to 11

OpenStudy (fibonaccichick666):

yes, but the work was incorrect, which is what I look at

OpenStudy (fibonaccichick666):

but anyways, like I said it doesn't matter at this point

OpenStudy (freckles):

the work isn't inccorrect

OpenStudy (freckles):

if -x->5 then x->-5

OpenStudy (freckles):

so i replaced the x with -5

OpenStudy (freckles):

i could have replaced -x with 5 or x with -5

OpenStudy (fibonaccichick666):

yes, it is \[lim_{−x→5}(−x+6)=\color\red{−(−5)}+6=5+6=11\] The step in red is not correct, based off of the limit you exchange the entire -x and turn it into a positive 5. You do not insert a negative 5 into x.

OpenStudy (fibonaccichick666):

we just said the same thing essentially, so anyways thanks for the help, have a good night

OpenStudy (freckles):

no you are saying what I was incorrect and i'm telling you it isn't

OpenStudy (freckles):

but anyways I think I'm done with this

OpenStudy (freckles):

but if you need someone else to clarify what I have said then I will holler at @ganeshie8 maybe he will be able to prove it to you

ganeshie8 (ganeshie8):

\[ \begin{align}\lim_{h\to 0} \frac{f(x)-f(x-h)}{h} &= \lim_{t=-h\to 0} \frac{f(x)-f(x+t)}{-t}\\~ \\&= \lim_{t=-h\to 0} \frac{f(x+t)-f(t)}{t}\\~\\ & = f'(x) \end{align}\]

ganeshie8 (ganeshie8):

oh its already there above, i have nothing else to say here :O http://prntscr.com/52v507

ganeshie8 (ganeshie8):

@Kainui

OpenStudy (ikram002p):

use definition of slope :)